A perfectly competitive firm has a cost function of C=Q2+20Q+700 and the equilibrium price is birr 100. find

A. The profit maximizing level of output

B. The maximum level of profit 


1
Expert's answer
2022-02-16T12:46:33-0500

A) "C=Q^2+20Q+700"

"EP=MR=100 birr"

Differentiating cost function yields;

"C'=MC=2Q+20"

Profit maximization level of output =MR=MC

"100=2Q+20"

"Q=4UNITS"


B)The maximum level of profit;

"=TR-TC"

"TC=4^2+20(4)+700"

"=796birr" "=400birr"

"TR=100Q=100*4=400 Birr"

"Profit=400-796=-396birr"

Negative profit implies financial loss.